OpenStudy (anonymous):

A cotton candy holder is shaped like a cone. The height is 12 in., and the diameter is 12 in. What is the volume of the holder? Use 3.14 to approximate pi, and express your final answer to the nearest tenth.. ______in3

OpenStudy (anonymous):

please help ill meadal

OpenStudy (anonymous):

First, do you know how to find the volume of a cone?

OpenStudy (anonymous):

yea i think i have to do it right now @alina1234

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the volume of a cone? Can you tell me?

OpenStudy (anonymous):

is the answer 1809 ?

OpenStudy (anonymous):

@alina1234

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

the volume of a cone is V=πr^2*h*1/3

OpenStudy (anonymous):

Let's start by finding the radius. The diameter is 12, and divide it by 2 to get the radius

OpenStudy (anonymous):

soooo..... now what lol

OpenStudy (anonymous):

The radius is 6 and notice in the formula it asked you to square it. So what's 6^2 or 6 squared?

OpenStudy (anonymous):

3? do u have to divide that

OpenStudy (anonymous):

No, 6^2 is actually 36

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

so 36?

OpenStudy (anonymous):

Yes. Then we times 36 by 3.14 (π)

OpenStudy (anonymous):

What's 36*3.14?

OpenStudy (anonymous):

50.24

OpenStudy (anonymous):

Do you have a calculator that can check your answer?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

can u do it?

OpenStudy (anonymous):

Good thing I do! The answer is 113.04. So 36*3.14=113.04

OpenStudy (anonymous):

so 113.04?

OpenStudy (anonymous):

Now according to the formula V=πr^2 h*1/3. Do you know what we do next?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So we have already solved πr^2, we times the answer we got 113.04 by the height (h)

OpenStudy (anonymous):

Can you tell me what the height is?

OpenStudy (anonymous):

113.04?

OpenStudy (anonymous):

Look back at the original problem, what is the height

OpenStudy (anonymous):

113.04 is πr^2

OpenStudy (anonymous):

oh yeaa

OpenStudy (anonymous):

so now what

OpenStudy (anonymous):

"A cotton candy holder is shaped like a cone. The height is 12 in., and the diameter is 12 in." that was the original problem, the height is 12 in

OpenStudy (anonymous):

Now we times πr^2 by h.

OpenStudy (anonymous):

Remember earlier I told you that πr^2 is 113.04?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

We times 113.04 by the height which is 12. What do you get? 113.04*12=...

OpenStudy (anonymous):

ummm im getting long numbers and its hard when u dont have a calculator lol

OpenStudy (anonymous):

You get 113.04*12=1356.48

OpenStudy (anonymous):

But that's not the answer yet! The last thing we need to do is divide by 3.

OpenStudy (anonymous):

So the final answer is 452.16 in^3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you're welcome